# Exploiting Minority Pseudo-Labels for Semi-Supervised Semantic Segmentation in Autonomous Driving

### v2

@@ Abstract
With the advancement of autonomous driving, semantic segmentation has achieved remarkable progress. The training of such networks heavily relies on image annotations, which are very expensive to obtain. Semi-supervised learning can utilize both labeled data and unlabeled data with the help of pseudo-labels. However, in many real-world scenarios where classes are imbalanced, majority classes often play a dominant role during training and the learning quality of minority classes can be undermined. To overcome this limitation, we propose a synergistic training framework, including a professional training module to enhance minority class learning and a general training module to learn more comprehensive semantic information. Based on a pixel selection strategy, they can iteratively learn from each other to reduce error accumulation and coupling. In addition, a dual contrastive learning with anchors is proposed to guarantee more distinct decision boundaries. In experiments, our framework demonstrates superior performance compared to state-of-the-art methods on benchmark datasets.
keywords:
Semi-supervised learning, Semantic segmentation, Autonomous driving
1 Introduction
Accurately segmenting different categories of road elements is crucial for effective semantic segmentation in intelligent transportation systems. Assigning a specific label to each pixel of pedestrians, vehicles, road signs, and other elements can enhance road safety and efficiency. The progress in this area can not only contribute to safer roadways but also pave the way for more intelligent and responsive driving systems. To this end, semantic segmentation becomes a critical task in computer vision, involves assigning a specific label to each pixel [1, 2, 3]. This pixel-level classification problem is notably more complex than image classification [4] or object detection [5] in the field of computer vision.
Many approaches to semantic segmentation utilize an encoder-decoder architecture [6, 7], a strategy often employed in fields such as defect detection and autonomous driving. The dependence of supervised learning techniques on large-scale, accurately annotated datasets, which are often costly to produce, has sparked considerable interest in leveraging weakly labeled, easily labeled, and unlabeled data to boost model performance [8, 9, 10, 11, 12, 13]. Specifically, there has been significant progress in semi-supervised semantic segmentation, which combines a small amount of labeled data with a large volume of unlabeled data.
@@ 1 Introduction
Semi-supervised semantic segmentation methods generally revolve around three core architectures: mean teacher [17, 14, 15], cross confidence consistency [16], and cross pseudo supervision [11]. In mean teacher architecture, the teacher model functions as an exponential moving average (EMA) of the student model [18]. The other two architectures involve two student models with the same architecture learning from each other. Additionally, two primary learning strategies are employed in semi-supervised semantic segmentation: consistency regularization [19, 20, 21] and entropy minimization [22, 23, 24]. Consistency regularization guarantees that a classifier generates similar predictions for a sample when subjected to various perturbations, and entropy minimization is a self-training process where a model is supervised using pseudo-labels generated by another model.
However, these approaches do not pay attention to class imbalance issues, and most of them are likely to suffer from model coupling issues [18, 25], both of which can lead to more severe error accumulation. To overcome these limitations, we propose a Synergistic Training framework with Professional and General Training (STPG). The simple yet effective framework consists of two modules: a professional training module and a general training module. The first module only selects pseudo-labels that are predicted consistently or highly mismatched between a professional student (Pro-Student) and a general teacher (Gen-Teacher), which can learn more minority class information and reduce error accumulation. In the second module, a general student (Gen-Student) is supervised by all pseudo-labels from a professional teacher (Pro-Teacher), which can learn more comprehensive semantic information. As the two modules are different, they are unlikely to be coupled. To our knowledge, this method represents a pioneering approach in semi-supervised semantic segmentation, addressing both class imbalance and model coupling issues simultaneously.
Our contributions are summarized as follows:
  • We propose a synergistic training framework with professional and general training modules, which promotes the learning of high-quality and minority-class pixels and reduces the influence of unreliable pseudo-labels.
  • To further improve the learning of minority classes, we propose dual contrastive learning with anchors which can highlight decision boundaries between different classes.
  • We demonstrate that our approach significantly boosts the model performance and outperforms the state-of-the-art techniques on two well-established datasets.
2 Related Work
@@
Semantic Segmentation plays a crucial role in computer vision, serving as a cornerstone for numerous applications. Recent advancements have witnessed remarkable strides in semantic segmentation, which are largely propelled by deep convolutional neural networks (CNNs) [26] and transformers [27]. The leading-edge approaches in semantic segmentation [28, 29, 30] predominantly adhere to the encoder-decoder paradigm [7]. In this paradigm, the decoder replaces fully connected layers that are typically employed in classification tasks with convolutional layers, enabling pixel-wise predictions. However, conventional supervised approaches heavily depend on large-scale, precisely annotated datasets, which can be prohibitively expensive to acquire in practice.
@@
Semi-Supervised Learning has seen significant advancements in recent years. Many of these methods share similar basic techniques, such as consistency regularization [31, 32, 33] or pseudo-labeling [34, 35]. Consistency regularization is based on the clustering assumption that a decision boundary typically passes through a region with low sample density; thus, predictions are consistent under perturbations. Pseudo-labeling is based on the natural idea that predictions obtained from a model can be reused as supervision, which is motivated by entropy minimization. For example, FixMatch [24] is notable for its technique of generating pseudo-labels from weakly augmented unlabeled images and utilizing them to supervise models fed with strongly augmented images. Long-tail distribution has been extensively studied in semi-supervised learning. DST [36] enhances the quality of pseudo-labels by adversarially optimizing representations to avoid worst-case scenarios. BaCon [37] employs a meticulously crafted contrastive approach to directly regularize the distribution of instance representations.
@@
Semi-Supervised Semantic Segmentation aims to utilize both unlabeled data and labeled data to train a high-performance segmentation network that can assign pseudo-labels to images. Due to limited labeled data, the full utilization of unlabeled data has become an important issue. Preliminary works have focused on consistency regularization and entropy minimization. U2PL [15] uses unreliable data for contrastive learning to boost performance. PRCL [38] introduces a framework that enhances representation quality by incorporating its probability. MMFA [39] introduces a framework for feature augmentation that incorporates both multi-reliability and multi-level strategies to fully exploit pixel information. CTT [25] alleviates model coupling issues through dual mean teacher architecture. Recently, numerous contrastive learning methods [40, 41] have been proposed for semi-supervised semantic segmentation. However, the above semi-supervised semantic segmentation methods do not sufficiently explore the semantic information within minority classes. In this manuscript, we propose a synergistic training framework with two training modules and a dual contrastive learning to resolve class imbalance issues.
3 Method
Figure 1: Overview of our framework. For labeled images, we apply weak augmentation for a labeled image and then feed it into Gen-Student and Pro-Student, supervised by the ground truth. For unlabeled images, we use two different modules: (a) In the professional training module, we apply weak augmentation and strong augmentation to an unlabeled image, feeding them into Gen-Teacher and Pro-Student, respectively, and then use refined pseudo-labels by pixel selection from Gen-Teacher to supervise Pro-Student’s prediction. (b) In the general training module, weak and strong augmentations are applied to an unlabeled image, which are then fed into Pro-Teacher and Gen-Student, respectively, followed by the utilization of all pseudo-labels from Pro-Teacher to supervise Gen-Student’s prediction. (c) In addition, we introduce a dual contrastive learning to foster distinct decision boundaries in the model, ensuring that it does not solely cater to the majority classes. (d) We evaluate the performance of STPG using Gen-Student.
@@ 3 Method
In this section, first, we define the problem. Second, we introduce the synergistic training framework. Last, we describe the utilization of dual contrastive learning with anchors.
3.1 Problem Definition
@@
3.2 Synergistic Training
Overview. To mitigate the accumulation of erroneous knowledge during training and reduce coupling, we employ a dual mean teacher architecture alongside mutual learning. In this setup, the parameters of the teacher model in one module are updated through EMA using the student model in other module. Fig. 1 summarizes the overall framework and the training procedure. This framework consists of two parallel student models, Gen-Student $f_{\theta_{Gen}}$ and Pro-Student $f_{\theta_{Pro}}$, each with different initializations. Their corresponding teacher models are Gen-Teacher $f_{\xi_{Gen}}$ and Pro-Teacher $f_{\xi_{Pro}}$. Specifically, ${f}_{\theta_{Pro}}$ and ${f}_{\xi_{Gen}}$ form a professional training module, and ${f}_{\theta_{Gen}}$ and ${f}_{\xi_{Pro}}$ form a general training module, each focusing on a specific aspect. With the professional training module, Pro-Student and Pro-Teacher can acquire more information from high-quality samples or from minority classes. In contrast, we hope that the general training module can learn more comprehensively.
@@ 3.2 Synergistic Training
Professional Training Module. The traditional learning approach tends to focus on majority classes and is not able to sufficiently learn minority classes. Thus, we propose a pixel selection strategy for producing refined pseudo-labels to improve the performance. The loss is calculated using consistent and highly mismatched pseudo-labels from Gen-Teacher to supervise Pro-Student’s predictions instead of all the pseudo-labels. The minority-class pseudo-labels are approximated with highly mismatched pseudo-labels, based on the observation that minority classes are more likely to be misclassified as other classes.
@@
Figure 2: Illustration of pixel selection strategy. $\mathbf{A}(\cdot)[i,j]$ represents the predicted class for the pixel located at the coordinates $(i,j)$. (a) When the two predictions of a pixel are consistent, they are used to train Pro-Student. (b) When the two predictions for a pixel are inconsistent, if the mismatch score of the class predicted by Gen-Teacher is larger than that of Pro-Student, the pixel usually contains more minority-class information and can be used for training Pro-Student. (c) Otherwise, these pseudo-labels are not used for training Pro-Student.
In this work, the pseudo-label represented by matrix $\hat{y}_{u}^{Gen}$ are divided into three mutually exclusive parts: consistent pseudo-labels, highly mismatched pseudo-labels and low mismatched pseudo-labels, represented by $\hat{y}_{u}^{Cons}$, $\hat{y}_{u}^{Hmis}$, $\hat{y}_{u}^{Lmis}\in\{0,1\}^{W\times H\times C}$. For each pixel along the third axis, only one of the three matrices can have a nonzero entry; namely, we can apply the following restriction: $\hat{y}_{u}^{Cons}+\hat{y}_{u}^{Hmis}+\hat{y}_{u}^{Lmis}=\hat{y}_{u}^{Gen}$.
Consistent pseudo-labels $\hat{y}_{u}^{Cons}$, as shown in Fig. 2(a), are defined as the pseudo-labels that both Gen-Teacher and Pro-Student have consistently predicted. Since the pixel predictions remain consistent across different training modules, we consider these pseudo-labels to be of high quality.
To identity mismatch pseudo-labels, we introduce a class-level mismatch score, which represents the proportion of a class being misclassified into other classes, based on the predictions of the Gen-teacher and Pro-student. This score reflects the degree of mismatch for a specific class. The highly mismatched pseudo-labels represented by $\hat{y}_{u}^{Hmis}$, as shown in Fig. 2(b), are computed in three steps.
(1) We construct the confusion matrix $M\in{\mathbb{R}^{C\times{C}}}$ with $C$ as the number of classes for the predictions of Pro-Student and Gen-Teacher in each mini-batch. An illustration is shown in Fig. 3.
@@
(2) We define a class-level mismatch score vector $I$ by the confusion matrix, in which each element represents the proportion of mismatched predictions of Pro-Student and Gen-Teacher for a certain class. For example, we compute the mismatch score of class $q$ as follows:
$I_{q}=\frac{{\textstyle\sum_{k=1}^{C}}m_{q,k}-m_{q,q}}{{\textstyle\sum_{k=1}^{C}m_{q,k}}}+\frac{{\textstyle\sum_{k=1}^{C}}m_{k,q}-m_{q,q}}{{\textstyle\sum_{k=1}^{C}m_{k,q}}}.$ (5)
When the score is high, it indicates significant disagreement between the two modules regarding Class $q$. Such substantial disagreement suggests that Class $q$ is likely a minority class, as these classes tend to be misclassified more often, resulting in fewer pixels being assigned to them in the pseudo-labels compared to other classes.
We map the class-level mismatch scores onto the pixel predictions of both the Pro-Student and Gen-Teacher to generate two mismatch matrices, denoted as $\mathbf{I}_{(\cdot)}$. These matrices match the size of the image, with each element corresponding to the mismatch score of the predicted class, namely, $\mathbf{I}_{(\cdot)}[i,j]=I_{\mathbf{A}(\cdot)[i,j]}$. Here, $\mathbf{A}(\cdot)[i,j]$ represents the predicted class for the pixel located at the coordinates $(i,j)$ and $\cdot$ denotes $p_{u}^{Pro}$ or $\hat{p}_{u}^{Gen}$
(3) For inconsistent predictions between Pro-Student and Gen-Teacher, we select certain pseudo-labels to form highly mismatched pseudo-labels:
$\hat{y}^{Hmis}_{u}=\hat{y}^{Gen}_{u}\times(\mathbf{I}_{(p_{u}^{Pro})}<\mathbf{I}_{(\hat{p}_{u}^{Gen})}),$ (6)
where $\mathbf{I}_{(p_{u}^{Pro})}<\mathbf{I}_{(\hat{p}_{u}^{Gen})}$ is a binary matrix with each entry equal to one only if the mismatch score of the Gen-Teacher’s prediction is greater than the Pro-Student’s prediction for the corresponding pixel. For example, for a certain pixel, if the Pro-Student’s prediction is Class $p$ and the Gen-Teacher’s prediction is Class $q$, and if $I_{p}<I_{q}$, we consider the Gen-Teacher’s prediction to be a highly mismatched pixel.
The low mismatched pseudo-labels $\hat{y}^{Lmis}_{u}$, as shown in Fig. 2(c), are the remaining pseudo-labels that do not belong to $\hat{y}^{Cons}_{u}$ or $\hat{y}^{Lmis}_{u}$. These pseudo-labels are not used for training Pro-Student.
@@
$\mathcal{L}_{u}^{Pro}=\omega_{u}^{Pro}\ell_{ce}(p_{u}^{Pro},\hat{y}_{u}^{Cons}+\hat{y}_{u}^{Hmis}),$ (7)
where $\omega_{u}^{Pro}$ is a weighting matrix and each element of $\omega_{u}^{Pro}$ represents the confidence of a pseudo-label generated by the teacher model, i.e., $\omega_{u}^{Pro}[i,j]=\max_{c\in\{1,...,C\}}\hat{p}_{u}^{Gen}[i,j,c]$. For pseudo-labels not belonging to $\hat{y}_{u}^{Cons}$ or $\hat{y}_{u}^{Hmis}$, the corresponding elements in $\omega_{u}^{Pro}$ are zero. Using the prediction confidence of the teacher model as the weight of the student model to calculate the loss can reduce the impact of suspicious noisy labels. Therefore, unlike traditional confidence thresholding approaches for filtering out numerous pseudo-labels, we can take advantage of unlabeled data.
@@
3.3 Dual Contrastive Learning with Anchors

### v3

# Exploiting Minority Pseudo-Labels for Semi-Supervised Fine-grained Road Scene Understanding

@@ Abstract
In fine-grained road scene understanding, semantic segmentation plays a crucial role in enabling vehicles to perceive and comprehend their surroundings. By assigning a specific class label to each pixel in an image, it allows for precise identification and localization of detailed road features, which is vital for high-quality scene understanding and downstream perception tasks. A key challenge in this domain lies in improving the recognition performance of minority classes while mitigating the dominance of majority classes, which is essential for achieving balanced and robust overall performance. However, traditional semi-supervised learning methods often train models overlooking the imbalance between classes. To address this issue, firstly, we propose a general training module that learns from all the pseudo-labels without a conventional filtering strategy. Secondly, we propose a professional training module to learn specifically from reliable minority-class pseudo-labels identified by a novel mismatch score metric. The two modules are crossly supervised by each other so that it reduces model coupling which is essential for semi-supervised learning. During contrastive learning, to avoid the dominance of the majority classes in the feature space, we propose a strategy to assign evenly distributed anchors for different classes in the feature space. Experimental results on multiple public benchmarks show that our method surpasses traditional approaches in recognizing tail classes.
I Introduction
Fine-grained road scene understanding is essential for intelligent transportation subsystems, including on-vehicle perception in autonomous cars, real-time traffic surveillance, and edge-based infrastructure monitoring, where accurate scene interpretation is critical to safety and performance. A fundamental technique to achieve this is semantic segmentation, which assigns a specific class label to each pixel in a road scene image, thereby identifying elements such as traffic signs, vehicles, pedestrians, cyclists, and other relevant obstacles [32, 10, 63]. Given its role in enabling such detailed scene comprehension, a robust semantic segmentation method can significantly improve situational awareness and support accurate decision-making, making it indispensable for developing safe and efficient intelligent transportation systems.
@@ I Introduction
Yet, supervised semantic segmentation for understanding large complex scenes is often costly due to the expensive annotation cost, and has sparked considerable interest in leveraging weakly labeled, easily labeled, and unlabeled data to boost model performance [48, 77, 73, 11, 9, 43]. Among them, semi-supervised semantic segmentation utilizes a small amount of labeled data with a large volume of unlabeled data to tackle this issue, generally revolve around three core architectures: mean teacher [54, 28, 60], cross confidence consistency [34], and cross pseudo supervision [11].In mean teacher architecture, the teacher model functions as an Exponential Moving Average (EMA) of the student model [35]. The other two architectures involve two student models with the same architecture learning from each other. Additionally, two primary learning strategies are employed in semi-supervised semantic segmentation: consistency regularization [3, 23, 50] and entropy minimization [2, 41, 52]. Consistency regularization guarantees that a classifier generates similar predictions for a sample when subjected to various perturbations, and entropy minimization is a self-training process where a model is supervised using pseudo-labels generated by another model.
Figure 1: Number of pixels in each class in Cityscapes dataset. The Cityscapes dataset exhibits a significant class imbalance, as reflected in the distribution of finely annotated pixels across different classes. The number of pixels per class, represented on the y-axis, varies widely, with certain dominant classes occupying the majority of the dataset’s pixel annotations, while other less prevalent classes account for only a small fraction.
However, in the aforementioned semi-supervised semantic segmentation approaches, there are still two critical challenges remaining. The first challenge lies in class imbalance, which is particularly pronounced in real-world road scene images. As shown in Fig. 1, pixel classes often exhibit long-tailed distributions, where majority classes (e.g., roads, sky) dominate the dataset, while safety-critical minority classes (e.g., traffic lights, riders) are significantly underrepresented. Traditional methods [6, 42, 61] are based on fixed prior knowledge for sampling or weighting, yet struggle to adapt to varying data characteristics[26]. To overcome this limitation, we introduce a cooperative training strategy consisting of a professional module that emphasizes minority-class pseudo-labels and a general module that learns from all pseudo-labels. This task-aware division improves flexibility while maintaining overall performance. In addition, to avoid the dominance of the majority classes in the feature space, we also introduce anchor contrastive learning such that different classes can be more evenly distributed in the embedded feature space by anchors.
The second challenge stems from model coupling. As training progresses, traditional mean-teacher models tend to become highly coupled [35, 62], causing them to learn highly similar semantic features. This limits their capacity to capture novel and useful information, ultimately hindering generalization. To alleviate this issue, we introduce a cross-training paradigm in which each student model is supervised by the teacher model from another module other than its own EMA [54] counterpart. By applying differentiated learning strategies, each module becomes semantically decoupled.
In short, to cater to traffic scenarios, we propose a Synergistic Training framework with Professional and General Training (STPG) by synergistically integrating professional and general training modules with cross-guided decoupling mechanisms, which can address both class imbalance and model coupling issues, and the contributions are summarized as follows:
  • We propose a synergistic training framework with a professional module that leverages a mismatch score metric to identify and learn from minority class pseudo-labels, working alongside a general module with cross-supervision to reduce model coupling.
  • We propose an anchor contrastive learning strategy using evenly distributed anchors and optimal anchor-prototype matching to alleviate majority-class feature dominance and thus enhance class discrimination.
  • Our method significantly improves recognition of minority classes without sacrificing the performance of majority classes, leading to more balanced and robust segmentation.
II Related Work
@@
Semantic Segmentation and Road Scene Understanding are essential for fine-grained perception of traffic environments, enabling accurate recognition of traffic signs, drivable areas, and other critical road elements, thereby forming the foundation for decision-making in intelligent transportation systems. Recent advancements have witnessed remarkable strides in semantic segmentation, which are largely propelled by deep convolutional neural networks (CNNs) [40] and transformers [55]. The leading-edge approaches in semantic segmentation [13, 24, 74] predominantly adhere to the encoder-decoder paradigm [45]. In this paradigm, the decoder replaces fully connected layers that are typically employed in classification tasks with convolutional layers, enabling pixel-wise predictions. For road scene understanding, SFNet-N[49] addresses boundary blurring in low-light images by integrating a light enhancement network with semantic information and a segmentation network with strong feature extraction capabilities. KGD[57] is a lightweight road segmentation framework for edge devices, where a student model learns from a teacher model using knowledge distillation to balance accuracy and efficiency. However, conventional supervised approaches heavily depend on large-scale, precisely annotated datasets, which can be prohibitively expensive to acquire in practice.
@@
Semi-Supervised Learning has seen significant advancements in recent years. Many of these methods share similar basic techniques, such as consistency regularization [17, 36, 7, 69] or pseudo-labeling [76, 30]. Consistency regularization is based on the clustering assumption that a decision boundary typically passes through a region with low sample density; thus, predictions are consistent under perturbations. Pseudo-labeling is based on the natural idea that predictions obtained from a model can be reused as supervision, which is motivated by entropy minimization. For example, FixMatch [52] is notable for its technique of generating pseudo-labels from weakly augmented unlabeled images and utilizing them to supervise models fed with strongly augmented images. Long-tail distribution has been extensively studied in semi-supervised learning. DST [8] enhances the quality of pseudo-labels by adversarially optimizing representations to avoid worst-case scenarios. BaCon [21] employs a meticulously crafted contrastive approach to directly regularize the distribution of instance representations.
@@
Semi-Supervised Semantic Segmentation aims to utilize both unlabeled data and labeled data to train a high-performance segmentation network that can assign pseudo-labels to images. Due to limited labeled data, the full utilization of unlabeled data has become an important issue. Preliminary works have focused on consistency regularization and entropy minimization. PRCL [64] introduces a framework that enhances representation quality by incorporating its probability. MMFA [70] introduces a framework for feature augmentation that incorporates both multi-reliability and multi-level strategies to fully exploit pixel information. UniMatch[65] proposes a two-stream perturbation technique so that two strong views are simultaneously guided by a common weak view. CorrMatch [53] introduces an adaptive threshold updating strategy and a pixel-wise similarity-based propagation of high-confidence predictions to efficiently leverage unlabeled data. However, the above semi-supervised semantic segmentation methods do not sufficiently explore the semantic information within minority classes. In this manuscript, we propose a synergistic training framework with two training modules and an anchor contrastive learning to resolve class imbalance issues.
Contrastive Learning is widely used to learn similarity functions by pulling representations of similar samples closer and pushing dissimilar ones farther apart in the representation space. There are many different variations of contrastive learning for task-specific purposes. CLSA[59] proposes a general framework that leverages the distribution divergence between weakly and strongly augmented images to enhance retrieval and representation learning. PaCo[15] proposes a parametric contrastive learning framework with class-wise learnable centers to address long-tailed recognition. DCL[67] addresses the inefficiencies in traditional contrastive learning by removing the negative-positive-coupling effect. Recently, numerous contrastive learning methods [1, 75, 12, 47] have been proposed for semi-supervised semantic segmentation. PPS [12] employs automatic clustering and contrastive learning to manage outliers and establish clear decision boundaries. ESL[47] proposes a pixel-to-part contrastive learning method cooperated with an unsupervised object part grouping mechanism to improve its ability to distinguish between different classes. Different from previous contrastive learning approaches, in this work, aiming at addressing the class imbalance issue, we introduce evenly distributed anchors in the feature space, which serve as reference points to improve class-level balance.
III Method
Figure 2: Overview of our framework. There are two training parts: professional and general, as illustrated in parts (a) and (b). In the professional training module, Pro-Student focusing on minority classes is supervised by selected pseudo-labels, and in the general training module, Gen-Student is supervised by all pseudo-labels. Teachers are updated Exponential Moving Average (EMA) [35] to maintain a smoothed version of past student model parameters. Anchor contrastive learning is designed to foster more evenly distributed decision boundaries when training labels from different classes are imbalanced. The supervised learning from labeled data is omitted for simplicity. The network architecture of each student or teacher model is illustrated in part (c).
@@ III Method
In this section, first, we define the problem. Second, we introduce the synergistic training framework. Last, we describe the utilization of anchor contrastive learning with anchors.
III-A Problem Definition
@@
III-B Synergistic Training
Overview. To mitigate the accumulation of erroneous knowledge during training and reduce coupling, we employ a dual mean teacher architecture alongside mutual learning. In this setup, the parameters of the teacher model in one module are updated through EMA using the student model in other module. Fig. 2 summarizes the overall framework and the training procedure, which consists of two parallel student models, a general student model (Gen-Student) $f_{\theta_{Gen}}$ and a professional student model (Pro-Student) $f_{\theta_{Pro}}$, each with different initializations. Their corresponding teacher models are general teacher model (Gen-Teacher) $f_{\xi_{Gen}}$ and professional teacher model (Pro-Teacher) $f_{\xi_{Pro}}$. Specifically, ${f}_{\theta_{Pro}}$ and ${f}_{\xi_{Gen}}$ form a professional training module, and ${f}_{\theta_{Gen}}$ and ${f}_{\xi_{Pro}}$ form a general training module, each focusing on a specific aspect. With the professional training module, Pro-Student and Pro-Teacher can acquire more information from high-quality samples or from minority classes. In contrast, we hope that the general training module can learn general feature representations from the whole dataset and the professional training module can focus on those from minority classes. Notably, all four models adhere to the same network architecture, which is based on DeepLabV3+[10]. The specific architecture is illustrated in the Fig. 2(c).
@@ III-B Synergistic Training
Professional Training Module. The traditional learning approach tends to focus on majority classes and is not able to sufficiently learn minority classes. Thus, we propose a pixel selection strategy for producing refined pseudo-labels to improve the performance. The loss is calculated using high-quality and high-mismatch pseudo-labels from Gen-Teacher to supervise Pro-Student’s predictions instead of all the pseudo-labels. The minority-class pseudo-labels are approximated with high-mismatch pseudo-labels, based on the observation that minority classes are more likely to be misclassified as other classes.
@@
Figure 3: Illustration of pixel selection strategy. $\mathbf{I}(\cdot)[i,j]$ represents the predicted class for the pixel located at the coordinates $(i,j)$. (a) When the two predictions of a pixel are consistent, they are considered high-quality and used to train Pro-Student. (b) When the two predictions for a pixel are inconsistent, if the mismatch score of the class predicted by Gen-Teacher is larger than that of Pro-Student, the pixel usually contains more minority-class information and can be used for training Pro-Student. (c) Otherwise, these pseudo-labels are not used for training Pro-Student.
In this work, the pseudo-labels represented by matrix $\hat{y}_{u}^{Gen}$ are divided into three mutually exclusive parts: high-quality pseudo-labels, high-mismatch pseudo-labels and low-mismatch pseudo-labels, represented by $\hat{y}_{u}^{Hq}$, $\hat{y}_{u}^{Hmis}$, $\hat{y}_{u}^{Lmis}\in\{0,1\}^{W\times H\times C}$. For each pixel along the third axis, only one of the three matrices can have a nonzero entry; namely, we can apply the following restriction: $\hat{y}_{u}^{Hq}+\hat{y}_{u}^{Hmis}+\hat{y}_{u}^{Lmis}=\hat{y}_{u}^{Gen}$.
(1) High-quality pseudo-labels $\hat{y}_{u}^{Hq}$, as shown in Fig. 3(a), are defined as the pseudo-labels that both Gen-Teacher and Pro-Student have consistently predicted. Inspired by DMT [22], we consider pseudo-labels with consistent pixel predictions across different training modules to be of high-quality. This is because the likelihood of both modules making the same error simultaneously is relatively low.
(2) To identity mismatch pseudo-labels, we introduce a class-level mismatch score, which represents the proportion of a class being misclassified into other classes, based on the predictions of the Gen-teacher and Pro-student. This score reflects the degree of mismatch for a specific class. The high-mismatch pseudo-labels represented by $\hat{y}_{u}^{Hmis}$, as shown in Fig. 3(b), are computed in three steps.
(i) We construct the confusion matrix $\mathbf{M}\in{\mathbb{R}^{C\times{C}}}$ with $C$ as the number of classes for the predictions of Pro-Student and Gen-Teacher in each mini-batch. An illustration is shown in Fig. 4.
@@
(ii) We define a class-level mismatch score vector $I$ by the confusion matrix, in which each element represents the proportion of mismatched predictions of Pro-Student and Gen-Teacher for a certain class. For example, we compute the mismatch score of class $q$ as follows:
$I_{q}=\frac{{\textstyle\sum_{k=1}^{C}}m_{q,k}-m_{q,q}}{{\textstyle\sum_{k=1}^{C}m_{q,k}}}+\frac{{\textstyle\sum_{k=1}^{C}}m_{k,q}-m_{q,q}}{{\textstyle\sum_{k=1}^{C}m_{k,q}}},$ (5)
where $m_{q,k}$ is the $(q,k)$-th entry of $\mathbf{M}$. Here, the row term ${{\textstyle(\sum_{k=1}^{C}}m_{q,k}-m_{q,q})}/{{\textstyle\sum_{k=1}^{C}m_{q,k}}}$ measures the proportion of pixels that Pro-Student predicts as class $q$ but Gen-Teacher disagrees with. The column term ${{\textstyle(\sum_{k=1}^{C}}m_{k,q}-m_{q,q})}/{{\textstyle\sum_{k=1}^{C}m_{k,q}}}$ quantifies the proportion of pixels that Gen-Teacher predicts as class $q$ but Pro-Student disagrees with. Thus, this combination offers a comprehensive measure of class-level disagreement: a high score reflects strong inconsistency between the two modules for Class $q$, which commonly occurs for minority or difficult classes. We therefore prioritize these classes, as they represent more challenging cases that require focused attention.
We map the class-level mismatch scores onto the pixel predictions of both the Pro-Student and Gen-Teacher to generate two mismatch matrices, denoted as $\mathbf{I}_{(p_{u}^{Pro})}$ and $\mathbf{I}_{(\hat{p}_{u}^{Gen})}$. These matrices match the size of the input image, where each element corresponds to the mismatch score of the predicted class:
$\mathbf{I}_{(\cdot)}[i,j]=I_{\arg\max_{k}(\cdot)[i,j,k]},$ (6)
Here, $I_{\arg\max_{k}(\cdot)[i,j,k]}$ is mismatch score associated with the predicted class for the pixel at coordinates $(i,j)$.
(iii) For inconsistent predictions between Pro-Student and Gen-Teacher, we select certain pseudo-labels to form high-mismatch pseudo-labels. We first compute a binary mask matrix $\mathcal{M}_{Hmis}$, where each element is defined as:
$\mathcal{M}_{Hmis}[i,j]=\mathbb{I}\left[\mathbf{I}_{(p_{u}^{Pro})}[i,j]<\mathbf{I}_{(\hat{p}_{u}^{Gen})}[i,j]\right],$ (7)
where $\mathbb{I}$ denotes indicator function. The inequality is performed element-wise, meaning that each element of $\mathcal{M}_{Hmis}$ is set to $1$ if the mismatch score of the Gen-Teacher’s predicted class is greater than that of the Pro-Student’s one for the corresponding pixel, and 0 otherwise.
Using this mask, we generate the high-mismatch pseudo-labels $\hat{y}^{Hmis}_{u}$ by retaining only the valid pixels from the Gen-Teacher’s predictions. Specifically, a pixel is considered valid if the corresponding element in the mask $\mathcal{M}_{Hmis}$ is 1, and invalid if the corresponding element is 0:
$\hat{y}^{Hmis}_{u}=\hat{y}^{Gen}_{u}\odot\mathcal{M}_{Hmis}.$ (8)
For example, for a certain pixel, if the Pro-Student’s prediction is Class $p$ and the Gen-Teacher’s prediction is Class $q$, and if $I_{p}<I_{q}$, we consider the Gen-Teacher’s prediction to be a high-mismatch pseudo-label. These pseudo-labels have high mismatch scores, indicating that they are easily misled by other majority classes. As a result, the model fails to learn these classes effectively. Therefore, we encourage the model to fully explore the semantic information of these classes to enhance their learning.
(3) The low-mismatch pseudo-labels $\hat{y}^{Lmis}_{u}$, as shown in Fig. 3(c), are the remaining pseudo-labels that do not belong to $\hat{y}^{Hq}_{u}$ or $\hat{y}^{Lmis}_{u}$. These pseudo-labels are not used for training Pro-Student.
@@
$\mathcal{L}_{u}^{Pro}=\omega_{u}^{Pro}\ell_{ce}(p_{u}^{Pro},\hat{y}_{u}^{Hq}+\hat{y}_{u}^{Hmis}),$ (9)
where $\omega_{u}^{Pro}$ is a weighting matrix and each element of $\omega_{u}^{Pro}$ represents the confidence of a pseudo-label generated by the teacher model, i.e., $\omega_{u}^{Pro}[i,j]=\max_{c\in\{1,...,C\}}\hat{p}_{u}^{Gen}[i,j,c]$. For pseudo-labels not belonging to $\hat{y}_{u}^{Hq}$ or $\hat{y}_{u}^{Hmis}$, the corresponding elements in $\omega_{u}^{Pro}$ are zero. Using the prediction confidence of the teacher model as the weight of the student model to calculate the loss can reduce the impact of suspicious noisy labels. Therefore, unlike traditional confidence thresholding approaches for filtering out numerous pseudo-labels, we can take advantage of unlabeled data.
Specifically, the high-quality pseudo-labels, high-mismatch pseudo-labels, and low-mismatch pseudo-labels generation rely only on simple matrix operations such as comparison, indexing, and element-wise addition. The time complexity for computing the confusion matrix in Step (i) involves $\mathcal{O}(N)$ operations, as each pixel is processed to update the matrix. Similarly, the mismatch score calculation in Step (ii) requires $\mathcal{O}(C^{2})$ operations, as it involves summing over rows and columns for each class. However, the operations for matching matrices and corresponding masks in Step (iii) are dominated by $\mathcal{O}(N)$. Given that $N\gg C$, the overall time and space complexity remains linear with respect to the number of pseudo-labels, i.e., the number of pixels.
@@
III-C Anchor Contrastive Learning